In the Ireland, 10% of the population is red haired.  Assuming mating is random, what is the frequency of heterozygotes in Ireland?
Answer #1

It is given that 10% of the individual in a population have red hair. The frequency of red hair individual is 0.10

The allelic frequency for red (r) allele is √0.10= 0.316227766

The allelic frequency of R is 1- r= 1-0.31622776=-l 0.683772234

The heterozygous frequency is 2× R×r = 2×0.316227766× 0.683772234= 0.432455532

On round off upt9 two decimal place, the answer is 0.43